• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Nathan

    Log entry time 23:59:39 on January15,2002

    Entry number 80349

    keyword=shift summary

    1. As we start the swing shift, the target group is cooling down the LH2 target. The HRS is now energized to 1.711 GeV/c and is parked at 33.97 deg. The calorimeter is in the 2A position at 46 deg, 12.0 m. During the day shift, 5.0 cm of polyethylene was installed at the front of the veto plane.

    2. At approx 2100, MCC starts to put pulsed beam into the hall.

    3. Controlled access requested at 2125 to check out bad vacuum in scattering chamber. At the end of the shift, the target gurus are still trying to find the problem.